# Data Trade and Consumer Privacy

### v2

@@ A.1 Proof of Lemma 1
The basic idea is to merge the signals that induce the same posterior (Bergemann and Morris, 2016). Let $\hat{\pi}(l|\theta)=\sum\limits_{s:\mu_{s}<v_{L}/v_{H}}\pi(s|\theta)$ and $\hat{\pi}(h|\theta)=\sum\limits_{s:\mu_{s}\geq v_{L}/v_{H}}\pi(s|\theta)$. Then, the data trade is captured by ($(\{l,h\},\hat{\pi})$. In the content after Lemma 1, we use the notation of ($(\{l,h\},\pi)$ for the binary direct data trade ($(\{l,h\},\hat{\pi})$ without confusions. ∎
A.2 Proof of Lemma 2
For a general service space $\mathcal{X}$, we have results in the following Lemma 3. Lemma 2 directly follows Lemma 3 with $\mathcal{X}=\{L,H\}$.
Lemma 3.
@@ A.2 Proof of Lemma 2
Proof.
To make the proof easy to read, we write out repeatedly the intermediary’s problem in the general space $\mathcal{X}$. That is, $\max\limits_{x,f,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$ subject to
@@
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
@@
because changes in $-\left(x_{H}-H\right)^{2}$ and $-f_{H}$ are canceled out, and $-2\delta\left(x_{H}-H\right)$ decreases. So $IC_{L}$ (14) is not affected. Case 2: when $L<x_{H}<H$, decreasing $\left(x_{H}-H\right)^{2}$ means increasing $x_{H}$, which will increase $\left(x_{H}-L\right)^{2}$. Then $IC_{L}$ (14) obviously holds in this case.
@@
because changes in $-\left(x_{L}-L\right)^{2}$ and $-f_{L}$ are canceled out, and $-2\delta\left(x_{L}-L\right)$ decreases as $x_{L}$ decreases. So increasing $f_{L}$ while decreasing $\left(x_{L}-L\right)^{2}$ by equal size will not affect $IC_{H}$ (12) and thus not affect all the constraints but increase the intermediary’s payoff. Hence, $x_{L}>L$ cannot be optimal.
In particular, with $\mathcal{X}=\{H,L\}$, $x_{L}=L$ and $x_{H}=H$ which proves Lemma 2.
A.3 Proof of Proposition 1
Proof.
When $\mu_{0}>\frac{v_{L}}{v_{H}}$. From $OB_{H}$ (8) and $OB_{L}$ (9)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
The obedience constraint for the producer to charge a high price (8) is redundant, we can ignore (8). The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$ as $p_{0}=v_{H}$. So the intermediary’s problem is simplified into
@@ A.3 Proof of Proposition 1
$\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)\leq\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)$ ($OB_{L}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ ($FE$)
@@
Hence, there are two cases following the above Lemma: $(1)$ only $IR_{L}$ is binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show Proposition 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
@@
$-\delta^{2}\leq f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
@@
$-\delta^{2}\leq\pi\left(l|H\right)\Delta v$
@@
A.4 Proof of Proposition 2
@@
When $\mu_{0}<\frac{v_{L}}{v_{H}}$. The obedience constraint $OB_{L}$ (9) is redundant. This is because $OB_{H}$ (8) and $OB_{L}$ (9) are respectively
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
and $\frac{v_{L}}{v_{H}}-\mu_{0}>0$. The intermediary’s problem is simplified into
@@
Note that at least one of $IR_{L}$ and $IR_{H}$ (18) is binding. Otherwise, increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$, $IC_{L}$, $IR_{H}$ (18), $IR_{L}$ and will increase the intermediary’s payoff. In addition, the following Lemma excludes the case of only $IR_{L}$ being binding.
Lemma 5.
@@
Given $IR_{L}$ is binding $f_{L}=V$, we need to show (18) is binding. Suppose (18) is not binding. Then $IC_{H}$ must be binding. Otherwise, increasing $f_{H}$ will not affect $IC_{H}$, (18), $IC_{L}$ and increase the intermediary’s payoff. So $IC_{H}$ is binding and it gives $f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$. From (18)
@@
Hence, there are two cases: $(1)$ both $IR_{L}$ and $IR_{H}$ (18) are binding; $(2)$ only $IR_{H}$ (18) is binding. Now we are ready to prove Proposition 2.
@@
In this step, we show that both $IR_{L}$ and $IR_{H}$ (18) are binding. Suppose not, then only $IR_{H}$ (18) is binding: $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$. As $IR_{L}$ is not binding, $IC_{L}$ must be binding. Otherwise, increasing $f_{L}$ will not affect $IR_{L}$, $IC_{L}$, $IC_{H}$ and will increase the intermediary’s payoff. So $IC_{L}$ is binding. The binding $IC_{L}$ and $IR_{H}$ (18) give $f_{L}=f_{H}+\delta^{2}=V+\left[\pi\left(l|H\right)-1\right]\Delta v+\delta^{2}$. From $IC_{H}$
$-\delta^{2}\leq\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
@@
When $\frac{\delta^{2}}{\Delta v}<1$, $\pi\left(l|H\right)\in\big{[}1-\frac{\delta^{2}}{\Delta v},1\big{]}$ and the payoff is $V$.
A.5 Proof of Corollary 3
@@
Suppose $\mu_{0}>\frac{v_{L}}{v_{H}}$ (the other case is similar). To see the full information data trade is optimal if and only if $\frac{\delta^{2}}{\Delta v}\geq 1$. Suppose the data trade is fully revealing $\pi(l|L)=1-\pi(l|H)=1$, $OB_{L}$ is satisfied in this situation. The maximum revenue that the intermediary can achieve under the full revelation data trade is the value of the following problem
$\max_{f_{L},f_{H}}$ $\quad f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+\left(1-\mu_{0}\right)v_{L}$
s.t. $f_{H}-f_{L}\leq\delta^{2}-\Delta v$
$f_{H}-f_{L}\geq-\delta^{2}$
$f_{H}\leq V$
$f_{L}\leq V$
where the constraints on service fees are from (12) (13) (14) (15). To make the choice set nonempty, it requires $\delta^{2}-\Delta v\geq-\delta^{2}$; otherwise, there is no solution to the problem, i.e., the full revelation data trade is not incentive compatible. The solution of $f_{H}$, $f_{L}$ and the maximum revenue are
  1. (i)
    for $\delta^{2}\geq\Delta v$, then $f_{H}=f_{L}=V$ and the revenue is $V+\left(1-\mu_{0}\right)v_{L}$. The full reveling achieves the maximum payoff for $\frac{\delta^{2}}{\Delta v}\geq 1$.
  2. (ii)
    for $\delta^{2}<\Delta v$, then $f_{L}=V$ and $f_{H}=V+\delta^{2}-\Delta v$ and the revenue is $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}\left(\delta^{2}-\Delta v\right)$. The full reveling is not optimal because the intermediary’s total revenue from full revelation is lower than that from the equilibrium one
    $V+\left(1-\mu_{0}\right)v_{L}+\mu_{0}\left(\delta^{2}-\Delta v\right)-[V+\left(1-\mu_{0}\right)v_{L}\frac{\delta^{2}}{\Delta v}]$
    $=$ $(\delta^{2}-\Delta v)\frac{\mu_{0}v_{H}-v_{L}}{\Delta v}<0$
    when $\delta^{2}<\Delta v$ and $\mu_{0}>\frac{v_{L}}{v_{H}}$.
A.6 Proof of Proposition 3
@@
Without a data market, the intermediary maximizes its service revenue by choosing the service provisions and fees
@@
s.t. $V-\left(x_{H}-H\right)^{2}-f_{H}\geq V-\left(x_{L}-H\right)^{2}-f_{L}$
@@
With the assumption $\mathcal{X}=\left\{L,H\right\}$, $x_{L}=L$ and $x_{H}=H$ by Lemma 3. Then the constraints turn into
$-f_{H}$ $\geq-\delta^{2}-f_{L}$
$-f_{L}$ $\geq-\delta^{2}-f_{H}$
$f_{H}$ $\leq V$
$f_{L}$ $\leq V$
Hence, $f_{L}=f_{H}=V$.
A.7 Proof of Proposition 5
@@
When $\mu_{0}>\frac{v_{L}}{v_{H}}$. From $OB_{H}$ (8) and $OB_{L}$ (9)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$$\geq$ 00 0
The obedience constraint for the producer to charge a high price $OB_{H}$ (8) is redundant, we can ignore (8). The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$. So the intermediary’s problem is simplified into

### v3

@@ A.1 Proof of 1
[…]
A.2 Proof of 2
For a general service space $\mathcal{X}=\mathbb{R}$, we have results in the following A.1. 2 directly follows A.1 with $\mathcal{X}=\{L,H\}$.
Lemma A.1.
@@ A.2 Proof of 2
Proof of A.1.
To make the proof easy to read, we write out explicitly the intermediary’s problem below. That is, $\max\limits_{x,f,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$ subject to
@@
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
@@
because changes in $\left(x_{H}-H\right)^{2}$ and $f_{H}$ are canceled out, and $2\delta\left(x_{H}-H\right)$ decreases. So $IC_{L}$ (13) is not affected. Case 2: when $L<x_{H}<H$, decreasing $\left(x_{H}-H\right)^{2}$ means increasing $x_{H}$, which will increase $\left(x_{H}-L\right)^{2}$. Then $IC_{L}$ (13) obviously holds in this case.
@@
because changes in $\left(x_{L}-L\right)^{2}$ and $f_{L}$ are canceled out, and $2\delta\left(x_{L}-L\right)$ decreases as $x_{L}$ decreases. So increasing $f_{L}$ while decreasing $\left(x_{L}-L\right)^{2}$ by equal size will not affect $IC_{H}$ (11) and thus not affect all the constraints but increase the intermediary’s payoff. Hence, $x_{L}>L$ cannot be optimal.
In particular, with $\mathcal{X}=\{H,L\}$, $x_{L}=L$ and $x_{H}=H$ which proves 2.
A.3 Proof of 1
Recall the obedience constraints in (6) and (7) are
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$ $\geq$ $\frac{v_{L}}{v_{H}}-\mu_{0}$ (18)
$\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)-\left(1-\frac{v_{L}}{v_{H}}\right)\mu_{0}\pi\left(l|H\right)$ $\geq$ $0$
where the second inequality is shortened by (16).
A.3.1 Proof of Part 1 of 1: $\mu_{0}>v_{L}/v_{H}$
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the obedience constraint (18) is redundant, we can ignore (18) and consider (16) only. The type-H consumers’ outside option $\max\left\{v_{H}-p_{0},0\right\}=0$ as $p_{0}=v_{H}$. So the intermediary’s problem is simplified into
@@ A.3.1 Proof of Part 1 of 1: $\mu_{0}>v_{L}/v_{H}$
$(1-\frac{v_{L}}{v_{H}})\mu_{0}\pi\left(l|H\right)\leq\frac{v_{L}}{v_{H}}\left(1-\mu_{0}\right)\pi\left(l|L\right)$ ($OB_{L}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
@@
Hence, there are two cases following the above Lemma: $(1)$ only $IR_{L}$ is binding; $(2)$ both $IR_{L}$ and $IR_{H}$ are binding. Now we show 1 by three steps: Step 1 shows that it’s not optimal that only $IR_{L}$ is binding; Step 2 simplifies the intermediary’s problem with binding $IR_{L}$ and $IR_{H}$; Step 3 solves the problem.
@@
$\delta^{2}\leq f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
@@
$\delta^{2}\leq\pi\left(l|H\right)\Delta v$
@@
A.3.2 Proof of Part 2 of 1: $\mu_{0}<v_{L}/v_{H}$
The detailed results of the part 2 of 1 are described below: Suppose $\mu_{0}<v_{L}/v_{H}$. In the optimal mechanism, the intermediary sets service fees to extract all the consumers surplus in the service and product markets, $f_{L}=V$, $f_{H}=V-\left(1-\pi\left(l|H\right)\right)\Delta v$ and the optimal data trade depends on the cross-market comparison between the mismatch cost $\delta^{2}$ and the vertical differentiation $\Delta v$: $\pi\left(l|L\right)=1$, $\pi\left(l|H\right)\in[\max\{1-\frac{\delta^{2}}{\Delta v},0\},1]$ The data trade fee is $T=\mu_{0}\Delta v\left(1-\pi\left(l|H\right)\right)$. And social welfare is $V+v_{L}+\mu_{0}\Delta v$.
@@ A.3.2 Proof of Part 2 of 1: $\mu_{0}<v_{L}/v_{H}$
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, the obedience constraint for the producer to charge a low price (16) is redundant and we consider (18) only. Moreover, the participation constraint of the type-H consumers (12) is
Without data trade the producer will charge $v_{L}$ and all consumers purchase in the product market. The producer’s revenue is $R_{0}=v_{L}$ without data trade. The net gain in the product revenue from data trade in (10) becomes
The intermediary’s problem is simplified into
@@
Note that at least one of $IR_{L}$ (14) and $IR_{H}$ (19) is binding. Otherwise, increasing $f_{L}$ and $f_{H}$ simultaneously by equal size will not affect $IC_{H}$ (11), $IC_{L}$ (13), $IR_{H}$ (19), $IR_{L}$ (14), and will increase the intermediary’s payoff. In addition, the following Lemma excludes the case of only $IR_{L}$ being binding.
Lemma A.3.
@@
Given $IR_{L}$ is binding $f_{L}=V$, we need to show (19) is binding. Suppose (19) is not binding. Then $IC_{H}$ must be binding. Otherwise, increasing $f_{H}$ will not affect $IC_{H}$, $IR_{H}$ (19), $IC_{L}$ and increase the intermediary’s payoff. So $IC_{H}$ is binding and it gives $f_{H}-V=\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$. From (19)
@@
Hence, there are two cases: $(1)$ both $IR_{L}$ and $IR_{H}$ (19) are binding; $(2)$ only $IR_{H}$ (19) is binding.
@@
In this step, we show that both $IR_{L}$ and $IR_{H}$ (19) are binding. Suppose not, then only $IR_{H}$ (19) is binding: $f_{H}=V+\left[\pi\left(l|H\right)-1\right]\Delta v$. As $IR_{L}$ is not binding, $IC_{L}$ must be binding. Otherwise, increasing $f_{L}$ will not affect $IR_{L}$ (14), $IC_{L}$ (13), $IC_{H}$ (11), and will increase the intermediary’s payoff. So $IC_{L}$ is binding. The binding $IC_{L}$ and $IR_{H}$ (19) give $f_{L}=f_{H}+\delta^{2}=V+\left[\pi\left(l|H\right)-1\right]\Delta v+\delta^{2}$. From $IC_{H}$
$\delta^{2}\leq\delta^{2}-\left[\pi\left(l|L\right)-\pi\left(l|H\right)\right]\Delta v$
@@
When $\frac{\delta^{2}}{\Delta v}<1$, $\pi\left(l|H\right)\in\big[1-\frac{\delta^{2}}{\Delta v},1\big]$ and the intermediary’s payoff is $V$.
A.4 Proof of 2
@@
Consider $\mu_{0}>\frac{v_{L}}{v_{H}}$ and $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}$ (the other case is similar). Suppose the data trade is of full information, $\pi(l|L)=1$ and $\pi(l|H)=0$. Then $OB_{L}$ holds. By plugging $f_{L}$ and $f_{H}$ into the objective, the intermediary chooses $x_{H}$, $x_{L}$, $\pi\left(l|L\right)$, $\pi\left(l|H\right)$ to maximize $V-\mu_{0}\left(x_{H}-H\right)^{2}-\left(1-\mu_{0}\right)\left(x_{L}-L\right)^{2}+\left(1-\mu_{0}\right)v_{L}$ subject to the following $IC_{L}$, $IC_{H}$
$\left(x_{H}-H\right)^{2}-\left(x_{H}-L\right)^{2}$ $\leq 0$ ($IC_{L}$)
$\left(x_{L}-L\right)^{2}-\left(x_{L}-H\right)^{2}$ $\leq-\Delta v$ ($IC_{H}$ )
So $x_{H}=H$, and if $\delta^{2}\geq\Delta v$, then $x_{L}=L$; if $\delta^{2}<\Delta v$, then $x_{L}=\frac{L+H}{2}-\frac{\Delta v}{2\delta}$. So given $\frac{\delta^{2}}{\Delta v}<\frac{\Delta v}{v_{L}+v_{H}}$, we have $x_{L}=\frac{L+H}{2}-\frac{\Delta v}{2\delta}$ and the intermediary’s payoff of selling full information is
$V-\left(1-\mu_{0}\right)\left(\frac{\delta}{2}-\frac{\Delta v}{2\delta}\right)^{2}+\left(1-\mu_{0}\right)v_{L}$
where the service profit is $V-\left(1-\mu_{0}\right)\left(\frac{\delta}{2}-\frac{\Delta v}{2\delta}\right)^{2}$ and the data profit is $\left(1-\mu_{0}\right)v_{L}$.
Recall that the intermediary’s payoff of the optimal mechanism is
$V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}$
So the difference between the two payoffs is positive
$V+(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}-\left[V-\left(1-\mu_{0}\right)\left(\frac{\delta}{2}-\frac{\Delta v}{2\delta}\right)^{2}+\left(1-\mu_{0}\right)v_{L}\right]$
$=$ $(1-\mu_{0})\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\left(1-\mu_{0}\right)\left(\frac{\delta}{2}-\frac{\Delta v}{2\delta}\right)^{2}-\left(1-\mu_{0}\right)v_{L}$
$\propto$ $\left(\frac{\delta}{\Delta v}\right)^{2}v_{L}v_{H}+\left(\frac{\delta}{2}-\frac{\Delta v}{2\delta}\right)^{2}-v_{L}$
$\propto$ $[(v_{L}+v_{H})\delta^{2}-(\Delta v)^{2}]^{2}>0$
which implies selling full information is not optimal.
A.5 Proof of 4
@@
Without a data market, the intermediary maximizes its service revenue by choosing the service provisions and fees to solve
@@
subject to $V-\left(x_{H}-H\right)^{2}-f_{H}\geq V-\left(x_{L}-H\right)^{2}-f_{L}$
@@
Under the assumption $\mathcal{X}=\left\{L,H\right\}$, $x_{L}=L$ and $x_{H}=H$ by A.1. Then the above constraints turn into $f_{H}\geq-\delta^{2}-f_{L}$, $f_{L}\geq-\delta^{2}-f_{H}$, $f_{H}\leq V$, $f_{L}\leq V$. Hence, $f_{L}=f_{H}=V$.
A.6 Proof of 2
@@
When $\mu_{0}<\frac{v_{L}}{v_{H}}$, the obedience constraint $OB_{L}$ is redundant and the type-H consumers’ outside option max $\left\{v_{H}-p_{0},0\right\}=v_{H}-v_{L}$. So the intermediary’s problem is simplified into
$\max_{f,x,\pi,T}f_{H}\mu_{0}+f_{L}\left(1-\mu_{0}\right)+T$
subject to the following constraints
$V-\left(x_{L}-L\right)^{2}-f_{L}\geq 0$ ($IR_{L}$)
$R\left(\pi\right)-R_{0}\geq T$ ($IR_{S}$)
$\pi\left(l|L\right),\pi\left(l|H\right)\in\left[0,1\right]$ (FEFE)
Lemma A.4.
If $IR_{L}$ (14) is binding, then $IR_{H}$ (19) is also binding.
Proof.
Given $IR_{L}$ is binding $f_{L}=V-\left(x_{L}-L\right)^{2}$. First, at least one of $IC_{H}$ and $IR_{H}$ is binding. Suppose not, both are not binding. Increasing $f_{H}$ will not affect $IC_{H}$, $IR_{H}$, $IC_{L}$ and increase the intermediary’s payoff.
Together with $f_{L}=V-\left(x_{L}-L\right)^{2}$, we have
where the inequality follows from $IR_{H}$. So $2x_{L}\geq L+H,$ leading a contradiction with $x_{L}\leq L.$ Hence, $IC_{H}$ is not binding.
Thus, by the above two points, $IR_{H}$ is binding. ∎
Step 1.
$f_{H}$ $=$ $V-\left(x_{H}-H\right)^{2}+(\pi\left(l|H\right)-1)\Delta v$
Then $IC_{H}$ is
$2\left(x_{L}-x_{H}\right)\leq[\pi\left(l|H\right)-\pi\left(l|L\right)]\frac{\Delta v}{\delta}$
and ignore (non-binding) $IR_{L}$ is
$2x_{H}-L-H<[1-\pi\left(l|H\right)]\frac{\Delta v}{\delta}$
The intermediary’s problem is
subject to $OB_{H}$ and $IC_{H}$: $2\left(x_{L}-x_{H}\right)\leq[\pi\left(l|H\right)-\pi\left(l|L\right)]\frac{\Delta v}{\delta}$, and $\pi\left(l|L\right)$, $\pi\left(l|H\right)\in\left[0,1\right]$. If $IC_{H}$ is not binding and ignore the constraint, the solution is
$2\left(x_{L}-x_{H}\right)=[\pi\left(l|H\right)-\pi\left(l|L\right)]\frac{\Delta v}{\delta}.$ (20)
Plugging (20) into the intermediary’s objective function, it maximizes
$V-\left(x_{H}-H\right)^{2}+\pi\left(l|H\right)\left(1-\mu_{0}\right)\Delta v+\left(1-\mu_{0}\right)\pi\left(l|L\right)v_{L}$
$+\Big[\left(x_{H}-L\right)^{2}-\Big(\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}\frac{\Delta v}{\delta}+x_{H}-L\Big)^{2}\Big]\left(1-\mu_{0}\right)-\left(1-\mu_{0}\right)v_{H}$
subject to $OB_{H}$, $\pi\left(l|L\right)\in\left[0,1\right]$, $\pi\left(l|H\right)\in\left[0,1\right]$. The first order derivative with respect to $\pi(l|H)$ is positive, so $OB_{H}$ is binding: $\pi(l|H)=\frac{\frac{v_{L}}{v_{H}}(1-\mu_{0})\pi(l|L)-(\frac{v_{L}}{v_{H}}-\mu_{0})}{(1-\frac{v_{L}}{v_{H}})\mu_{0}}$.
$x_{H}-L$ $=$ $\frac{\delta v_{L}}{v_{L}-\mu_{0}v_{H}}-\frac{\pi(l|H)-\pi(l|L)}{2}\frac{\Delta v}{\delta}$
$x_{H}-H$ $=$ $-\left(1-\mu_{0}\right)\frac{\pi\left(l|H\right)-\pi\left(l|L\right)}{2}\frac{\Delta v}{\delta}$
from which $H-(1-\mu_{0})L-\frac{(1-\mu_{0})\delta v_{L}}{v_{L}-\mu_{0}v_{H}}=\mu_{0}x_{H}$ and then
$\mu_{0}(x_{H}-H)$ $=$ $H-(1-\mu_{0})L-\frac{(1-\mu_{0})\delta v_{L}}{v_{L}-\mu_{0}v_{H}}-\mu_{0}H$
$=$ $(1-\mu_{0})\delta\frac{-\mu_{0}v_{H}}{v_{L}-\mu_{0}v_{H}}<0$
which contradicts with $x_{H}\geq H$. Hence, $\pi(l|L)=1$ is binding: $\pi(l|L)=1$. And $\pi(l|H)=1$ by the binding $OB_{H}$, then $x_{H}=H$. By (20), $x_{L}=x_{H}=H$ which is a contradiction with $x_{L}\leq L$. Hence, it’s not feasible that only $IR_{H}$ is binding.
Step 2.
Following Step 1, it must be both $IR_{L}$ and $IR_{H}$ are binding:
$f_{L}$ $=$ $V-\left(x_{L}-L\right)^{2}$
Then from $IC_{H}$
and from $IC_{L}$
equivalently, they are
A.7 Proof of 3
Proof.
When $\mu_{0}>\frac{v_{L}}{v_{H}}$, the obedience constraint $OB_{H}$ (18) is redundant, we ignore (18) and consider only (16) for obedience constraints. The type-H consumers’ outside option $\max\{v_{H}-p_{0},0\}=0$. So the intermediary’s problem is simplified into